Copper that was 63% pure was melted together with copper that was 90% pure to make 18 kilograms of an alloy that was 75% pure. How many kilograms of each kind were used?
Answer by stanbon(75887)

Copper that was 63% pure was melted together with copper that was 90% pure to make 18 kilograms of an alloy that was 75% pure. How many kilograms of each kind were used?
------
Equation:
copper + copper = copper
---
0.63x + 0.90(18-x) = 0.75*18
----
63x + 90*18 - 90x = 75*18
-27x = -15*18
---
x = 10 kg (amt of 63% alloy needed)
---
18-x = 8 kg (amt of 90% alloy needed)
==============
Cheers,
Stan H.